# Does a Rising Tide Lift All Boats? Bias Mitigation for AI-based CMR Segmentation

### v1

@@ 3 Results
The results of the baseline and the three bias mitigation methods can be seen in Table 3 and in Fig. 1. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median performance for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other races compared to the baseline, as can be seen in Fig. 1(a) and Fig. 1(b). This is perhaps surprising as performance on these other races was not optimised during training.
The other mitigation methods did not significantly improve performance for the Black subjects. Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing. Group DRO resulted in increased performance for the Black subjects but performance remained significantly lower than for the White subjects. Both oversampling and Group DRO slightly decreased median DSC for White subjects (although not statistically significantly), but reweighing significantly decreased median DSC for White subjects.
Table 3: DSC values for each of the bias mitigation methods. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
Baseline
Oversampling
Reweighing
Group DRO
@@
As shown in Table 4, the baseline model trained using cropped images improved performance for both White and Black subjects and reduced bias. Oversampling significantly improved performance for the Black subjects compared to the baseline, resulting in performance that was higher than for White subjects. Group DRO improved the DSC for Black subjects but this increase was not significant. Reweighing made performance slightly worse for both races, with median DSC decreasing and SER and FG measures increasing.
@@
Table 4: DSC values for each of the bias mitigation methods using cropped images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
@@
Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median 0.919 0.893 0.913 0.922 0.916 0.887 0.918 0.908 IQR 0.032 0.041 0.030 0.039 0.033 0.043 0.033 0.038 SER 1.326 1.115 1.352 1.130 Fairness gap 0.026 --0.009 0.029 0.011
@@
Finally, combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
@@
Table 5: DSC values for the combined bias mitigation methods using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
@@
Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
@@ 4 Discussion
This work has performed a comprehensive examination of bias mitigation methods for AI segmentation models used for cine CMR images. We have shown that bias in CMR segmentation models can be mitigated by using such methods. In particular, oversampling minority subjects reduces bias so that there is no significant difference between the performance of the Black and White subjects. Although oversampling did not add any extra information to the dataset, the method allowed the network to train on Black subjects more frequently than if oversampling was not used, allowing for better balance between protected groups. It could be anticipated that training using a small number of (oversampled) Black subjects would increase the risk of overfitting to those subjects, having a detrimental effect on generalisation, but this effect was not seen in our experiments as test performance remained high. Previous work in [10] showed that race could be classified from cine CMR images, indicating that there are distinct features in the images of different races that are recognisable to AI models. Using oversampling will allow the network to see more of these distinct features to learn better representations of the under-represented group.
Reweighing did not improve segmentation performance, instead decreasing performance for both protected groups. This may be due to increased importance being given to a small group of subjects, decreasing focus on the larger group. As described in [2] and [1], when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set. This suggests that segmentation of the White subjects’ images may be a more difficult task as there may be more outliers and variation in the hearts than in the Black subjects. Reweighing these White subjects so that their importance is lower in the loss function could decrease accuracy.
Using cropped images increased performance and reduced bias compared to using uncropped images. Using oversampling with cropped images reduced bias further, but interestingly not to the same extent as using uncropped images. Also, note that to use a model trained using cropped images in clinical practice, a region-of-interest proposal network would be required to identify the cropping region at test time [15].
@@
This work has some limitations. For example, only Black and White subjects were used during training to control for other factors which may affect mitigation methods. Future work could investigate the effect of mitigating biases using multiple races or in other protected attributes such as age and socioeconomic status. It would also be useful to investigate the performance of bias mitigation strategies under external validation.
5 Acknowledgements
@@
This work was supported by the Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1. This research has been conducted using the UK Biobank Resource under Application Number 17806

### v2

@@ 5 Results
5.1 Internal validation
The internal validation results of the baseline and bias mitigation methods can be seen in Table 4 and Fig. 2. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median DSC for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other ethnicities compared to the baseline (see Fig. 2). This is perhaps surprising as performance on these other ethnicities was not optimised during training.
@@ 5.1 Internal validation
The other mitigation methods did not significantly improve performance for the Black subjects. Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing and HD increasing. Group DRO resulted in increased performance for the Black subjects but performance remained significantly lower than for the White subjects. Both oversampling and Group DRO slightly decreased median DSC for White subjects (although not statistically significantly), but reweighing significantly decreased median DSC for White subjects.
@@
Table 4: DSC and HD values for each of the bias mitigation methods tested on the internal validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR DSC 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 Median HD (mm) 6.725 9.364 6.835 7.164 7.174 10.037 6.802 8.286 IQR HD (mm) 3.707 4.725 3.713 3.861 4.207 4.902 3.581 4.157 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
@@
Combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
@@
Table 5: DSC and HD values for the combined bias mitigation methods tested on the internal validation set using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median DSC 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR DSC 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 Median HD (mm) 6.725 9.364 6.985 7.740 6.831 8.417 7.364 8.654 IQR HD (mm) 3.707 4.725 3.666 4.079 3.520 4.001 4.320 4.558 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
As shown in Table 6, the baseline model trained using cropped images improved performance for both White and Black subjects and reduced bias. Oversampling significantly improved performance for the Black subjects compared to the baseline, resulting in performance that was higher than for White subjects. Group DRO improved the DSC for Black subjects compared to the baseline but this increase was not significant. However, the HD for both groups of subjects decreased. Reweighing made performance slightly worse for both ethnicities, with median DSC decreasing and SER and FG measures increasing.
Table 6: DSC and HD values for each of the bias mitigation methods tested on the internal validation set using cropped images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.917 0.888 0.906 0.914 0.915 0.877 0.917 0.903 IQR DSC 0.039 0.047 0.031 0.044 0.036 0.052 0.035 0.046 Median HD (mm) 5.713 7.895 6.394 6.479 5.987 8.945 5.700 7.170 IQR HD (mm) 2.669 3.960 3.359 3.211 3.073 4.097 2.805 3.584 SER 1.340 1.088 1.448 1.161 Fairness gap 0.028 -0.008 0.038 0.013
Table 7 and Figs. 2(e) to 2(f) show the results for the cascaded cropping on the internal validation dataset. Surprisingly, using the cascaded cropping approach significantly improved the DSC score compared to using the ground-truth cropped results shown in Table 6 for all bias mitigation methods. We discuss a possible reason for this in Section 6.1. Using oversampling on the cascaded cropping approach did not improve DSC score but decreased the HD, SER and FG compared to the baseline. Reweighing also improved the SER and FG.
Table 7: DSC and HD values for each of the bias mitigation methods using the cascaded cropping approach on the internal validation data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.959 0.937 0.958 0.941 0.951 0.935 0.955 0.943 IQR DSC 0.060 0.057 0.064 0.058 0.067 0.070 0.068 0.058 Median HD (mm) 3.757 6.564 3.731 4.929 5.176 6.426 3.841 4.623 IQR HD (mm) 2.268 5.012 2.129 2.717 3.262 4.172 2.039 3.327 SER 1.517 1.380 1.332 1.276 Fairness gap 0.021 0.016 0.016 0.012
5.2 External validation
Table 8 shows the external validation results. Tables S1 to S3 show the results broken down by cardiac region. As oversampling was found to be the best bias mitigation on internal validation, only the results for this model and the baseline model are shown. Note that, as the LVM is not routinely segmented for the ES frame in clinical workflows at our institution, results for the external validation dataset are shown for the ED frames only.
Table 8: DSC and HD values for the baseline and oversampling methods tested on the external validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Baseline p = 0.093 Oversampling p = 0.16 White Black White Black Median DSC 0.886 0.879 0.885 0.872 IQR DSC 0.030 0.043 0.037 0.038 Median HD (mm) 5.728 5.913 6.152 5.919 IQR HD (mm) 1.522 1.966 1.238 1.840 SER 1.064 1.106 Fairness gap 0.007 0.012
Overall, the model had high performance on the external validation set. The performance was comparable to that of the bias mitigation methods tested on the internal validation set, with the baseline performance for the Black subjects being higher than for the Black subjects in the internal validation set. The fairness gap and SER were also smaller than for the internal validation set. However, for the external validation set, using oversampling decreased performance for both groups, with the FG, SER and median HD increasing.
The approach based on cropping using ground truth segmentations produced a significantly better baseline performance in the external validation set for both groups compared to not using cropping, as shown in Table 9. Using oversampling slightly decreased the performance for both groups and resulted in a significantly different performance for Black and White subjects.
The results from using the cascaded cropping based approach can be seen in Table 9. Tables S4 to S6 show the results broken down by cardiac region for the cascaded approach. Using this approach, performance was lower than the upper bound of using ground truth segmentations for cropping for both groups but performance remained higher than for the baseline model using uncropped images seen in Table 8. Performance increased for both groups when using oversampling and the FG decreased. Table S7 shows the results for the oversampling cascaded model broken down by sex and age.
Table 9: DSC and HD values for the models using cropping on the external validation dataset. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Cropped baseline ground-truth model p = 0.069 Cropped oversampling ground-truth model Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 White Black White Black White Black White Black Median DSC 0.931 0.916 0.918 0.906 0.917 0.903 0.928 0.920 IQR DSC 0.043 0.037 0.040 0.033 0.045 0.038 0.039 0.026 Median HD (mm) 4.087 4.090 4.793 4.594 4.772 4.954 4.658 3.875 IQR HD (mm) 1.630 1.178 1.321 1.529 3.124 1.919 2.69 1. 432 SER 1.207 1.141 1.168 1.118 Fairness gap 0.014 0.012 0.014 0.008
@@ 6 Discussion
6.1 Internal validation
This work has performed a comprehensive examination of bias mitigation methods for AI segmentation models used for cine CMR images. We have shown that bias in CMR segmentation models can be mitigated by using such methods. In particular, oversampling minority subjects reduces bias so that there is no significant difference between the performance of the Black and White subjects. Although oversampling did not add any extra information to the dataset, the method allowed the network to train on Black subjects more frequently than if oversampling was not used, allowing for better balance between protected groups. It could be anticipated that training using a small number of (oversampled) Black subjects would increase the risk of overfitting to those subjects, having a detrimental effect on generalisation, but this effect was not seen in our experiments as test performance remained high. Previous work in Lee et al. (2025a) has showed that ethnicity can be classified from cine CMR images, indicating that there are distinct features in the images of different ethnicities that are recognisable to AI models. Using oversampling will allow the network to see more of these distinct features to learn better representations for the under-represented group.
Reweighing did not improve segmentation performance, instead decreasing performance for both protected groups. This may be due to increased importance being given to a small group of subjects, decreasing focus on the larger group. As described in Lee et al. (2022) and Lee et al. (2023), when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set. This suggests that segmentation of the White subjects’ images may be a more difficult task as there may be more outliers and variation in the hearts than in the Black subjects. Reweighing these White subjects so that their importance is lower in the loss function could decrease accuracy.
@@ 6.1 Internal validation
Using cropped images increased performance and reduced bias compared to using uncropped images. Using oversampling with cropped images reduced bias further, but interestingly not to the same extent as using uncropped images. The cascaded cropping approach resulted in a performance that was higher than using ground truth segmentations for cropping. This may be due to the intermediate predictions on the uncropped images (c in Fig. 1) being less accurate and the segmented area being larger than the ground truth segmentations, resulting in a bounding box which was larger and contained more information. This suggests that performance could be optimised further by increasing the size of the pixel buffer used in the cropping process.
6.2 External validation
Using the bias mitigation models on the external validation set produced comparable DSC scores to internal validation results. No significant bias was observed in the baseline model for this dataset, as shown in Table 8. This is surprising as the dataset is out-of-distribution and some distributional shift is expected, which could lead to worse performance. However, unlike the models tested on the internal validation set, oversampling did not decrease the FG or improve the performance of the models under external validation. This is consistent with previous work that has reported limited effectiveness of generic bias mitigation algorithms under complex domain shifts (Schrouff et al., 2022; Yang et al., 2024).
Cropping the images increased performance overall, both using the ground truth-based cropping and the cascaded approach. Importantly, using the cascaded approach, performance was higher than the baseline model using uncropped images for both groups under external validation. Interestingly, using oversampling in combination with the cascaded approach improved performance for both groups further. These results indicate that, not only does cropping improve the performance of models for both protected groups (i.e. ethnicities), but also that ground truth segmentations are not needed to crop these images. This finding can be vital for the clinical translation of bias mitigation algorithms in CMR analysis, as better segmentation performance will allow for better assessment of clinical biomarkers and better treatment planning, prognosis and diagnosis. Provided that there is both a network trained on uncropped images and a network trained on cropped images, the method will be deployable and scalable. In the future, the first model used to localise the cardiac region of interest (b in Fig. 1) could be replaced with a similar network trained for bounding box detection such as that proposed in He et al. (2017).
The best method overall (cascaded cropping used with oversampling) was based on an understanding of the root cause of bias in AI-based CMR analysis as reported in Lee et al. (2025a). This method resulted in not only reduced bias compared to the baseline model but also improved performance for both groups. This important result highlights that, when using a bias understanding-informed mitigation approach, there was no fairness-accuracy trade-off. The fairness-accuracy trade-off is sometimes used as a reason not to implement bias mitigation algorithms in clinical practice, due to the medical principle of non-maleficence. Our work suggests that this trade-off is not inevitable, but that it can be avoided by a careful analysis of bias in individual scenarios rather than application of a generic mitigation algorithm.
6.3 Limitations and future work
@@
This work has some limitations. For example, only Black and White subjects were used during training. Future work could investigate the effect of mitigating biases using multiple ethnicities, in other protected attributes such as age and socioeconomic status, or in intersectional groups. There was also a relatively small number of subjects in the external dataset when compared to the number of subjects in the training and internal validation sets from the UK Biobank, so externally validating on a larger, more diverse dataset would be beneficial.
7 Conclusions
This paper has reported the most comprehensive investigation of bias mitigation in AI-based CMR segmentation to date. We have shown that the fairness-accuracy trade-off can be avoided by using a bias understanding-informed approach to mitigation, rather than using a generic mitigation algorithm. Therefore, this represents an important finding that should motivate further investigations into such bias understanding-informed approaches to mitigation in other applications.
[…]
@@ 7 Conclusions
Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1. This research has been conducted using the UK Biobank Resource under Application Number 17806.
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals and human subjects
[…]
E.P-A. is an employee of Heartflow Inc. All work presented is independent of her role at Heartfow. P-G.M. worked as a consultant for Perspectum Diagnostics Ltd until 2023. The remaining authors declare no conflict of interest.
[…]
The UK Biobank dataset is publicly available for approved research projects. Requests to access the dataset should be directed to https://www.ukbiobank.ac.uk/. The clinical dataset cannot be publicly shared due to limitations of the ethical approval.
Supplementary Material
Table S1: Metrics for LVBP segmentation based on inference performed on external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Baseline p = 0.11 Oversampling p = 0.46 White Black White Black Median DSC 0.940 0.931 0.936 0.937 IQR DSC 0.041 0.035 0.038 0.032 Median HD (mm) 3.929 3.416 3.705 3.518 IQR HD (mm) 1.622 1.498 1.296 2.186 SER 1.146 1.007 Fairness gap 0.009 -0.0004
Table S2: Metrics for LVM segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Baseline p = 0.84 Oversampling p = 0.57 White Black White Black Median DSC 0.817 0.803 0.808 0.802 IQR DSC 0.046 0.075 0.053 0.06 Median HD (mm) 5.026 5.141 5.096 5.318 IQR HD (mm) 1.583 1.941 1.995 2.042 SER 1.077 1.034 Fairness gap 0.014 0.006
Table S3: Metrics for RVBP segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Baseline p = 0.43 Oversampling p = 0.27 White Black White Black Median DSC 0.883 0.882 0.885 0.868 IQR DSC 0.057 0.048 0.070 0.071 Median HD (mm) 8.902 7.783 8.869 10.181 IQR HD (mm) 3.208 3.096 3.962 4.401 SER 1.004 1.148 Fairness gap 0.0004 0.017
Table S4: Metrics for LVBP segmentation based on inference performed on the external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Cropped baseline p = 0.088 Cropped oversampling p = 0.76 Cropped baseline cascaded model Cropped oversampling cascaded model p = 0.72 White Black White Black White Black White Black Median DSC 0.962 0.949 0.958 0.943 0.959 0.943 0.959 0.946 IQR DSC 0.030 0.031 0.024 0.027 0.031 0.029 0.036 0.024 Median HD (mm) 2.887 2.488 3.043 2.870 3.132 2.839 2.842 2.493 IQR HD (mm) 0.984 0.854 1.113 1.697 1.119 1.049 1.352 2.646 SER 1.309 1.364 1.398 1.316 Fairness gap 0.012 0.015 0.016 0.013
Table S5: Metrics for LVM segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Cropped baseline p = 0.39 Cropped oversampling p = 0.63 Cropped baseline cascaded model p = 0.69 Cropped oversampling cascaded model p = 0.86 White Black White Black White Black White Black Median DSC 0.881 0.860 0.852 0.849 0.856 0.855 0.879 0.863 IQR DSC 7 0.103 0.093 0.096 0.083 0.093 0.106 0.071 0.066 Median HD (mm) 3.512 3.310 4.071 3.937 4.051 4.008 3.422 3.259 IQR HD (mm) 1.738 1.251 2.030 2.492 2.532 2.098 2.051 3.120 SER 1.175 1.019 1.009 1.130 Fairness gap 0.021 0.003 0.001 0.016
Table S6: Metrics for RVBP segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
Cropped baseline p = 0.75 Cropped oversampling p = 0.71 Cropped baseline cascaded model p =0.84 Cropped oversampling cascaded model White Black White Black White Black White Black Median DSC 0.938 0.925 0.918 0.907 0.912 0.905 0.929 0.932 IQR DSC 0.024 0.022 0.034 0.026 0.045 0.040 0.049 0.031 Median HD (mm) 5.932 5.787 6.540 6.662 7.477 7.147 4.871 6.426 IQR HD (mm) 2.580 3.557 2.991 2.753 4.459 5.708 2.018 4.328 SER 1.197 1.136 1.131 1.036 Fairness gap 0.012 0.011 0.011 -0.003
Table S7: DSC and HD values by sex and age for the cascaded approach using oversampling on the external validation set. The p-values were computed based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. FG was calculated using $DSC_{male}-DSC_{female}$ and $DSC_{Age<50}-DSC_{Age>50}$
Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 Female Male Age $<$ 50 Age $>$ 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003